# Compositional Design of Safety Controllers for Large-scale Stochastic Hybrid Systems

### v1

@@ 2.4. Interconnected SHS
The interconnection of $\Phi_{i}$ is $\Phi=(X,U,\mathcal{U},\sigma,\rho,f_{1},\varsigma,f_{2})$, denoted by $\mathcal{I}(\Phi_{1},\ldots,\Phi_{N})$, such that $X=\Pi_{i=1}^{N}X_{i},U=\Pi_{i=1}^{N}U_{i},\sigma:=\mathsf{blkdiag}(\sigma_{1}(x_{1}),\ldots,\sigma_{N}(x_{N})),\rho:=\mathsf{blkdiag}(\rho_{1}(x_{1}),\ldots,\rho_{N}(x_{N})),f_{1}:=[f_{1_{1}};\ldots;f_{1_{N}}],f_{2}:=[f_{2_{1}};\ldots;\\ f_{2_{N}}]$ and $\varsigma:=[\varsigma_{1};\dots;\varsigma_{N}]$, subjected to the following interconnection constraint:
@@
$\Phi\!:\left\{\begin{array}[]{rl}\mathsf{d}x(t)=\!\!&f_{1}(x(t),\nu(t))\mathsf{d}t+\sigma(x(t))\mathsf{d}\mathbb{W}_{t}+\rho(x(t))\mathsf{d}\mathbb{P}_{t},~{}\quad\quad\quad~{}t\in\mathbb{R}_{0}^{+}\backslash\Delta,\\ x(t)=\!\!&f_{2}(x(t^{-}),\nu(t),\varsigma(t))\hskip 132.87431ptt\in\Delta.\end{array}\right.$ (2.9)
@@
An interconnection of two SHS $\Phi_{i}$ is schematically depicted in Fig. 2.
Figure 2. Interconnection of two SHS $\Phi_{1},\Phi_{2}$.
2.5. Augmented SHS
As it can be observed, the SHS in (2.3) encompasses both continuous and discrete evolutions, in which continuous behaviours are governed by stochastic differential equations with Brownian motions and Poisson processes, and instantaneous jumps are modelled by stochastic difference equations with additive noises. In the next definition, we provide a unified framework for SHS covering both continuous and discrete behaviours by adding an auxiliary variable $\vartheta_{i}$ to the state tuple of $\Phi_{i}$ as a counter, based on which an SHS is allowed or prevented from jump.
@@ 2.5. Augmented SHS
Consider an SHS $\Phi_{i}=(X_{i},U_{i},\mathcal{U}_{i},W_{i},\mathcal{W}_{i},\sigma_{i},\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$ with jump parameters ($\tau$, $\varepsilon_{1_{i}}$, $\varepsilon_{2_{i}}$). An augmented SHS (A-SHS) $\mathcal{A}_{i}(\Phi_{i})$ associated to $\Phi_{i}$ is characterized by $\mathcal{A}_{i}(\Phi_{i})\!=\!(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\\ \mathbb{Y}_{i},$ $\mathbb{H}_{i})$, where:
  • $\mathbb{X}_{i}=X_{i}\times\{0,\ldots,\varepsilon_{2_{i}}\}$ is the state space of A-SHS, while $(x_{i},\vartheta_{i})\in\mathbb{X}_{i}$ indicates that the current state is $x_{i}$, and the time elapsed since the latest jump is $\vartheta_{i}$;
@@
      Flow case: $10\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-$ 1, $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, and $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
@@
Note that $(x_{i}^{\prime},\vartheta_{i}^{\prime})=\mathbb{F}_{i}((x_{i},\vartheta_{i}),\nu_{i},w_{i},\sigma_{i},\rho_{i},\varsigma_{i})$ in Definition 2.3 characterizes the evolution of A-SHS, in which $\mathbb{F}_{i}$ is acquired based on one of the following two cases:
- if the counter $\vartheta_{i}$ is between 00 and $\varepsilon_{2_{i}}-1$, then the state evolves based on the flow case as $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, in which $x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$ is the solution of the stochastic differential equation $\mathsf{d}x_{i}(t)=f_{1_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t))\mathsf{d}t\!+\!\sigma_{i}(x_{i}(t))\mathsf{d}\mathbb{W}_{t}+\rho_{i}(x_{i}(t))\mathsf{d}\mathbb{P}_{t},$ which is well defined in the literature [ØS05]. In this case, the next counter is increased as $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
- if the counter $\vartheta_{i}$ is between $\varepsilon_{1_{i}}$ and $\varepsilon_{2_{i}}$, then the state evolves based on the jump case as $x_{i}^{\prime}=f_{2_{i}}(x_{i}(t^{-}),\nu_{i}(t),\\ w_{i}(t),\varsigma_{i}(t))$ and the counter is reset from zero as $\vartheta_{i}^{\prime}=0$.
In the following proposition, we demonstrate that the state trajectories of an A-SHS and its original SHS are equivalent. Consequently, we leverage the A-SHS as a unified framework that encompasses both continuous evolutions and instantaneous jumps, providing a more manageable model.
@@
3. Augmented Control (Sub-)Barrier Certificates
@@
Given an A-SHS $\mathcal{A}_{i}(\Phi_{i})=(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},\mathbb{H}_{i})$, consider $\mathbb{X}_{0_{i}}=X_{0_{i}}\times\{0\},~{}\mathbb{X}_{u_{i}}=X_{u_{i}}\times\{0,\ldots,\varepsilon_{2_{i}}\}$, as initial and unsafe regions of $\mathcal{A}_{i}(\Phi_{i})$, respectively, with $X_{0_{i}},X_{u_{i}}\subseteq X_{i}$ being initial and unsafe regions of $\Phi_{i}$. A function $\mathcal{B}_{i}:\mathbb{X}_{i}\to{\mathbb{R}}_{0}^{+}$ is said to be an augmented control sub-barrier certificate (A-CSBC) for $\mathcal{A}_{i}(\Phi_{i})$ if there exist $10<\gamma_{i}<$ 1, $\kappa_{i}\in{\mathbb{R}}^{+},\mu_{i},\beta_{i},\varphi_{i},\eta_{i}\in{\mathbb{R}}_{0}^{+}$, such that
$\mathcal{B}_{i}(x_{i},\vartheta_{i})\geq\kappa_{i}\|x_{i}\|^{2},\quad\quad\quad\quad\quad~{}\!\!\forall(x_{i},\vartheta_{i})\in\mathbb{X}_{i},$ (3.1a)
@@
We now provide a notion of augmented control barrier certificates for augmented SHS without disturbance signals, i.e., interconnected A-SHS (cf. Definition 4.1).
@@
Given an interconnected A-SHS $\mathcal{A}(\Phi)=(\mathbb{X},\mathbb{U},\sigma,\rho,\varsigma,\mathbb{F},\mathbb{Y},\mathbb{H})$ with initial and unsafe regions $\mathbb{X}_{0},\mathbb{X}_{u}$, a function $\mathcal{B}:\mathbb{X}\to{\mathbb{R}}_{0}^{+}$ is said to be an augmented control barrier certificate (A-CBC) for $\mathcal{A}(\Phi)$ if there exist $10<\gamma<$ 1, $\mu,\beta,\eta\in{\mathbb{R}}_{0}^{+}$, with $\beta>\mu$, such that
@@
Remark 3.4.
It is important to note that merely satisfying conditions (3.1)does not guarantee the safety of the network. In particular, an A-CSBC captures the influence of disturbance inputs $w_{i}$ through interaction gains $\varphi_{i}$ in  (3.1d), reflecting the impact of neighboring subsystems within the network topology. Under certain small-gain compositional conditions detailed in Section 4, these A-CSBCs are combined to form an A-CBC for the interconnected A-SHS, ensuring its safety. To do so, condition (3.1a) is crucial in deriving these compositional results (see (4.3)).
Note that $\beta>\mu$ is required in Definition 3.3 in order to ensure there is no intersection between initial and unsafe sets of interconnected A-SHS. The next theorem, borrowed from [Kus67] but adapted to A-SHS, shows the usefulness of A-CBC by quantifying a guaranteed probabilistic bound on the safety of the interconnected SHS.
@@
Given interconnected A-SHS $\mathcal{A}(\Phi)=(\mathbb{X},\mathbb{U},\sigma,\rho,\varsigma,\mathbb{F},\mathbb{Y},\mathbb{H})$, let $\mathcal{B}$ be an A-CBC for $\mathcal{A}(\Phi)$. Then for any $x_{0}$ and $\vartheta_{0}$, the probability that the state trajectories of $\mathcal{A}(\Phi)$ reach the unsafe region $\mathbb{X}_{u}$ within the finite-time horizon $k\in[0,\mathcal{T}]$ is computed as:
$\mathds{P}\Big{\{}x_{x_{0}\nu}(t_{k})\in\mathbb{X}_{u}~{}\text{for some}~{}k\in[0,\mathcal{T}]\,\,\big{|}\,\,x_{0},\vartheta_{0}\Big{\}}\leq\begin{cases}1-(1-\frac{\mu}{\beta})(1-\frac{\eta}{\beta})^{\mathcal{T}}\!,\quad\quad&\text{if }\beta\geq\frac{\eta}{1-{\gamma}},\\ (\frac{\mu}{\beta})\gamma^{\mathcal{T}}+(\frac{\eta}{(1-{\gamma})\beta})(1-\gamma^{\mathcal{T}}),\quad\quad&\text{if }\beta<\frac{\eta}{1-{\gamma}}.\\ \end{cases}$ (3.3)
@@
$\mathds{P}\Big{\{}x_{x_{0}\nu}(t_{k})\in\mathbb{X}_{u}~{}\text{for some}~{}k\in[0,\infty)\,\,\big{|}\,\,x_{0},\vartheta_{0}\Big{\}}\leq\frac{\mu}{\beta}.$ (3.4)
@@
$\mathds{P}$$\Big{\{}(x(t_{k}),\vartheta(t_{k}))\in\mathbb{X}_{u}\text{ for some}~{}0\leq k\leq\mathcal{T}\,\,\big{|}\,\,x_{0},\vartheta_{0}\Big{\}}\leq\mathds{P}\Big{\{}\sup_{0\leq k\leq\mathcal{T}}\mathcal{B}(x(t_{k}),\vartheta(t_{k}))\geq\beta\,\,\big{|}\,\,x_{0},\vartheta_{0}\Big{\}}.$ (3.5)
@@
4. Compositional Construction of A-CBC
In this section, we analyze networks of A-SHS via small-gain reasoning and discuss how to construct an A-CBC of the interconnected A-SHS based on the corresponding A-CSBCs of individual subsystems. To do so, we first introduce a formal definition for the interconnected A-SHS similar to Definition 2.2.

### v2

@@ 2.4 Interconnected SHS
The interconnection of $\Phi_{i}$ is $\Phi=(X,U,\mathcal{U},\sigma,\rho,f_{1},\varsigma,f_{2})$, denoted by $\mathcal{I}(\Phi_{1},\ldots,\Phi_{N})$, such that $X=\Pi_{i=1}^{N}X_{i},U=\Pi_{i=1}^{N}U_{i},\sigma:=\mathsf{blkdiag}(\sigma_{1}(x_{1}),\ldots,\sigma_{N}(x_{N})),\rho:=\mathsf{blkdiag}\\ (\rho_{1}(x_{1}),\ldots,\rho_{N}(x_{N})),f_{1}:=[f_{1_{1}};\ldots;f_{1_{N}}],f_{2}:=[f_{2_{1}};\ldots;f_{2_{N}}]$ and $\varsigma:=[\varsigma_{1};\dots;\varsigma_{N}]$, subjected to the following interconnection constraint:
@@
$\Phi\!:\left\{\hskip-4.2679pt\begin{array}[]{rl}\mathsf{d}x(t)=\!\!&f_{1}(x(t),\nu(t))\mathsf{d}t+\sigma(x(t))\mathsf{d}\mathbb{W}_{t}+\rho(x(t))\mathsf{d}\mathbb{P}_{t},\\ &~\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad~t\in\mathbb{R}_{0}^{+}\backslash\Delta,\\ x(t)=\!\!&{f_{2}(x(t^{-}),\nu(t),\varsigma(t))},\,\quad\quad\quad\quad\quad~\!\!t\in\Delta.\end{array}\right$. (2.11)
@@
We now formally define the safety specification of interest for the interconnected SHS.
@@
Given a safety specification $\Lambda=(X_{0},X_{u},\mathcal{T})$, with $X_{0},X_{u}\subseteq X$ representing the initial and unsafe sets of the SHS $\Phi$, where $X_{0}\cap X_{u}=\emptyset$, the system $\Phi$ is said to be safe within the time horizon $\mathcal{T}\in\mathbb{N}$, denoted by $\Phi\vDash\Lambda$, if all trajectories starting from $X_{0}$ avoid $X_{u}$ during $\mathcal{T}$. Since the trajectories of $\Phi$ are stochastic, we are interested in computing the probability $\mathds{P}\{\Phi\vDash\Lambda\}\geq 1-\epsilon$, for some $\epsilon\in(0,1]$.
2.5 Augmented SHS
As can be observed, the SHS in (2.4) encompasses both continuous and discrete evolutions, in which continuous behaviors are governed by stochastic differential equations with Brownian motions and Poisson processes, and instantaneous jumps are modeled by discrete-time stochastic reset equations with additive noises. To enable a unified treatment of these two modes of evolution—and to handle the logic governing jump timing—we introduce an augmented representation of SHS, adopted from (Swikir et al., 2020, Def. 5) but adapted to the stochastic setting, as formalized in the following definition.
Definition 2.8.
Consider an SHS $\Phi_{i}=(X_{i},U_{i},\mathcal{U}_{i},W_{i},\\ \mathcal{W}_{i},\sigma_{i},\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$ with jump parameters ($\tau$, $\varepsilon_{1_{i}}$, $\varepsilon_{2_{i}}$). An augmented SHS (A-SHS) $\mathcal{A}_{i}(\Phi_{i})$ associated to $\Phi_{i}$ is characterized by $\mathcal{A}_{i}(\Phi_{i})\!=\!(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},$ $\mathbb{H}_{i})$, where:
  • $\mathbb{X}_{i}=X_{i}\times\{0,\ldots,\varepsilon_{2_{i}}\}$ is the state space of A-SHS, where $(x_{i},\vartheta_{i})\in\mathbb{X}_{i}$ indicates that the current state is $x_{i}$, and the time elapsed since the latest jump is $\vartheta_{i}$;
@@ 2.5 Augmented SHS
      Flow case: $0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1$,
      $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, and $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
@@
      Jump case: $\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}}$,
      $x_{i}^{\prime}=f_{2_{i}}(x_{i}(\tau^{-}),\nu_{i}(\tau),w_{i}(\tau^{-}),\varsigma_{i}(\tau))$, and $\vartheta_{i}^{\prime}=0$;
@@
Note that $(x_{i}^{\prime},\vartheta_{i}^{\prime})=\mathbb{F}_{i}((x_{i},\vartheta_{i}),\nu_{i},w_{i},\sigma_{i},\\ \rho_{i},\varsigma_{i})$ in Definition 2.8 characterizes the evolution of A-SHS, in which $\mathbb{F}_{i}$ is acquired based on one of the following two cases:
  • If the counter $\vartheta_{i}$ is between 0 and $\varepsilon_{2_{i}}-1$, then the state may evolve based on the flow case as $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, in which $x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$ is the solution of the stochastic differential equation $\mathsf{d}x_{i}(t)=f_{1_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t))\mathsf{d}t\!+\!\sigma_{i}(x_{i}(t))\mathsf{d}\mathbb{W}_{t}+\rho_{i}(x_{i}(t))\mathsf{d}\mathbb{P}_{t},$ which is well defined in the literature (Øksendal & Sulem, 2005). In this case, the next counter is increased as $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
  • If the counter $\vartheta_{i}$ is between $\varepsilon_{1_{i}}$ and $\varepsilon_{2_{i}}$, then the state may evolve based on the jump case as $x_{i}^{\prime}={f_{2_{i}}(x_{i}(\tau^{-}),\nu_{i}(\tau),w_{i}(\tau^{-}),\varsigma_{i}(\tau))}$ and the counter is reset from zero as $\vartheta_{i}^{\prime}=0$.
We further clarify the evolution of the SHS and the A-SHS in Definitions 2.1 and 2.8, respectively, and how they are related to each other through the following illustrative example.
Example 2.10.
Consider the case where $\varepsilon_{1_{i}}=1$ and $\varepsilon_{2_{i}}=3$. Under Definition 2.1, the system evolves continuously according to the stochastic differential equation until a jump occurs. If the first jump happens at time $t_{k}$, then the next jump at $t_{k+1}$ must satisfy $t_{k+1}-t_{k}\in\{\tau,2\tau,3\tau\},$ meaning that successive jumps occur no sooner than $\tau$ and no later than $3\tau$ after the preceding jump. If no jump takes place at $\tau$ or $2\tau$ following $t_{k}$, a jump is guaranteed at $3\tau$. This pattern repeats for every pair of consecutive jumps.
In Definition 2.8, the auxiliary state variable $\vartheta_{i}$ acts as a counter functioning as memory to track the elapsed time between consecutive jumps, and it increments by one at each time step of length $\tau$. Hence, when $\varepsilon_{1_{i}}=1$ and $\varepsilon_{2_{i}}=3$, while $\vartheta_{i}$ (an integer value) remains within the inclusive range $\varepsilon_{1_{i}},\varepsilon_{2_{i}}]$, a jump may occur—specifically at $\vartheta_{i}=1$ (after $\tau$) or at $\vartheta_{i}=2$ (after $2\tau$). If no jump happens by the time $\vartheta_{i}$ reaches $\varepsilon_{2_{i}}=3$, a jump is then enforced at $3\tau$. Immediately following any jump, $\vartheta_{i}$ resets to zero, beginning a new count toward the next permissible jump interval. More precisely, the system begins in the flow case with $\vartheta_{i}=0$, after which the counter increments to one ($\vartheta_{i}=1$). Then, since $\varepsilon_{1_{i}}\leq\vartheta_{i}\leq\varepsilon_{2_{i}}$, a jump might happen, and if so, the system evolves following the jump case. Otherwise, $\vartheta_{i}$ increases to two ($\vartheta_{i}=2$). If a jump does not happen again, $\vartheta_{i}$ becomes three ($\vartheta_{i}=3$). Since $0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1$ does not hold anymore, the system necessarily evolves according to the jump case, and $\vartheta_{i}$ resets to zero to count the time elapsed until the next jump. Thereby, Definitions 2.1 and 2.8 represent the same model.
Although the equivalency of Definitions 2.1 and 2.8 is intuitively demonstrated thorough Example 2.10, in the following proposition, we formally show that the state trajectories of an A-SHS and its original SHS are equivalent. Consequently, we leverage the A-SHS as a unified framework that encompasses both continuous evolutions and instantaneous jumps, providing a more manageable model.
@@
3 Augmented Control (Sub-)Barrier Certificates
@@
The following definition formalizes the notion of an augmented control sub-barrier certificate (A-CSBC), which acts as a local certificate for each subsystem, borrowed from (Lavaei, Di Lillo, Censi & Frazzoli, 2022, Def. 3.3), but adapted to the SHS.
@@
Given an A-SHS $\mathcal{A}_{i}(\Phi_{i})=(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\\ \sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},\mathbb{H}_{i})$, consider $\mathbb{X}_{0_{i}}=X_{0_{i}}\times\{0\},~\mathbb{X}_{u_{i}}=X_{u_{i}}\times\{0,\ldots,\varepsilon_{2_{i}}\}$, as initial and unsafe regions of $\mathcal{A}_{i}(\Phi_{i})$, respectively, with $X_{0_{i}},X_{u_{i}}\subseteq X_{i}$ being initial and unsafe regions of $\Phi_{i}$. A function $\mathcal{B}_{i}:\mathbb{X}_{i}\to{\mathbb{R}}_{0}^{+}$ is said to be an augmented control sub-barrier certificate (A-CSBC) for $\mathcal{A}_{i}(\Phi_{i})$ if there exist $0<\gamma_{i}<1$, $\kappa_{i}\in{\mathbb{R}}^{+},\mu_{i},\beta_{i},\varphi_{i},\eta_{i}\in{\mathbb{R}}_{0}^{+}$, such that
$\mathcal{B}_{i}(x_{i},\vartheta_{i})\geq\kappa_{i}\|x_{i}\|^{2},\quad\quad\quad\quad~\!\forall(x_{i},\vartheta_{i})\in\mathbb{X}_{i},$ (3.1a)
@@
In this definition, the key inequality (3.1d) expresses how the expected certificate value evolves under one-step transitions. The presence of $\varphi_{i}$ reflects the influence of interconnections, which becomes crucial in compositional safety analysis.
@@
This remark highlights a key design advantage: the control input $\nu_{i}$ is independent of the internal coupling term $w_{i}$, thereby enabling fully decentralized controller synthesis. This structural property facilitates scalable control design without the need for global coordination among subsystems. We now provide a notion of augmented control barrier certificates for augmented SHS without disturbance signals, i.e., interconnected A-SHS (cf. Definition 4.1), adapted from (Lavaei, Di Lillo, Censi & Frazzoli, 2022, Def. 3.1).
@@
Given an interconnected A-SHS $\mathcal{A}(\Phi)=(\mathbb{X},\mathbb{U},\sigma,\rho,\varsigma,\mathbb{F},\mathbb{Y},\mathbb{H})$ with initial and unsafe regions $\mathbb{X}_{0},\mathbb{X}_{u}$, a function $\mathcal{B}:\mathbb{X}\to{\mathbb{R}}_{0}^{+}$ is said to be an augmented control barrier certificate (A-CBC) for $\mathcal{A}(\Phi)$ if there exist $0<\gamma<1$, $\mu,\beta,\eta\in{\mathbb{R}}_{0}^{+}$, with $\beta>\mu$, such that
@@
The A-CBC defined in Definition 3.2 serves as a global safety certificate, where conditions (3.2a)–(3.2b) ensure separation between initial and unsafe sets, while (3.2c) ensures expected safety evolution under controlled dynamics.
Note that $\beta>\mu$ is required in Definition 3.3 in order to ensure there is no intersection between initial and unsafe sets of interconnected A-SHS. The next theorem, borrowed from (Kushner, 1967) but adapted to A-SHS, shows the usefulness of A-CBC by quantifying a guaranteed probabilistic bound on the safety of the interconnected SHS, as a function of the certificate bounds and decay parameters.
@@
Given interconnected A-SHS $\mathcal{A}(\Phi)=(\mathbb{X},\mathbb{U},\sigma,\rho,\varsigma,\mathbb{F},\mathbb{Y},\mathbb{H})$, let $\mathcal{B}$ be an A-CBC for $\mathcal{A}(\Phi)$. Then for any $x_{0}$ and $\vartheta_{0}$, the probability that the state trajectories of $\mathcal{A}(\Phi)$ reach the unsafe region $\mathbb{X}_{u}$ within the finite-time horizon $k\in{[0,1,\dots\mathcal{T}]}$, is computed as:
$\mathds{P}\Big\{x_{x_{0}\nu}(t_{k})\in\mathbb{X}_{u}~\text{for some}~k\in{[0,1,\dots\mathcal{T}]}\,\,\big|\,\,x_{0},\vartheta_{0}\Big\}$
[…]
@@
$\mathds{P}\Big\{x_{x_{0}\nu}(t_{k})\!\in\!\mathbb{X}_{u}~\text{for some}~k\in[0,\infty)\,\,\big|\,\,x_{0},\vartheta_{0}\Big\}={\epsilon}\leq\frac{\mu}{\beta}.$ (3.4)
@@
$\mathds{P}$ $\Big\{(x(t_{k}),\vartheta(t_{k}))\in\mathbb{X}_{u}\text{ for some}~0\leq k\leq\mathcal{T}\,\,\big|\,\,x_{0},\vartheta_{0}\Big\}$
$\leq\mathds{P}\Big\{\sup_{0\leq k\leq\mathcal{T}}\mathcal{B}(x(t_{k}),\vartheta(t_{k}))\geq\beta\,\,\big|\,\,x_{0},\vartheta_{0}\Big\}.$ (3.5)
@@
4 Compositional Construction of A-CBC
Here, we first introduce a formal definition for the interconnected A-SHS similar to Definition 2.6.